# Sample-Efficiency in Multi-Batch Reinforcement Learning: The Need for Dimension-Dependent Adaptivity

### v1

@@ C.3.1 Preliminary Lemmas
This lemma is similar to Lemma C.4 but uses a more explicit notion of distance between subspaces. The proof is given in Appendix C.3.4.
C.3.2 Proof of Lemma C.3
To prove Lemma C.3, we show the existence of a subspace packing $\mathcal{C}$ in $\mathcal{G}_{m,d}(\mathbb{R})$ (with $m=2^{\lceil N/4\rceil}$) and use Lemma C.4. The subspace packing must satisfy two conditions:
@@
To show the existence of a suitable subspace packing, we use Lemma C.2 with $k=\lfloor\frac{g(\gamma)}{m}\sqrt{d}+1\rfloor$ ($\leq d$). Lemma C.2 gives a packing $\mathcal{C}$ of size $|\mathcal{C}|\geq\Big{(}\frac{d}{2}\Big{)}^{\frac{g(\gamma)}{2m}\sqrt{d}-1}\lfloor\frac{d}{m}\rfloor$ s.t
@@
the packing $\mathcal{C}$ satisfies $|\mathcal{C}|\geq n+1$ and $d_{\min}(\mathcal{C})\geq\sqrt{m-g(\gamma)^{2}}$. By Lemma C.4 there exists a subspace $A\in G_{m,d}(\mathbb{R})$ of dimension $m=2^{\lceil N/4\rceil}$ s.t
@@
C.3.3 Proof of Lemma C.4
Consider distinct $A,B\in\mathcal{C}$, i.e $A,B$ are subspaces of dimension $m$ s.t
@@
Now letting $20\leq\theta_{1}\leq...\leq\theta_{m}\leq\pi/$ 2 be the principal angles between $A$ and $B$ (see Appendix C), we have
@@
The first principal angle $\theta_{1}=\arccos a^{T}b$ where $a\in A$, $b\in B$ are unit vectors chosen s.t
@@
Since this applies to arbitrary distinct $A,B\in\mathcal{C}$, there are $(n+1)$ subspaces $A_{1},...,A_{n+1}\in\mathcal{C}$ of dimension $m$ s.t for any $i\neq j$,
@@
By Lemma C.5, there exists a subspace $H$ in $\{A_{1},...,A_{n+1}\}$ s.t for all $y\in\{y_{1},...,y_{n}\}$,
@@
C.3.4 Proof of Lemma C.5
Recall that we consider $(n+1)$ subspaces $A_{1},...,A_{n+1}$ of dimension $m$ s.t for any $i\neq j$,
@@
Appendix D Proof of Theorem 4.4
D.1 MDP Class
The PE problems used in the proof of Theorem 4.4 are characterised by a class of MDPs $\mathcal{M}$ and target policies $\Pi$. In this section, we define the class of MDPs $\mathcal{M}$. All MDPs $M\in\mathcal{M}$ in the class share the same state-space $\mathcal{S}$, action space $\mathcal{A}$, feature map and transition function $p$ but differ in the reward function $r_{M}$ and target policy $\pi_{M}$. This class of MDPs is the same as the one used by Zanette (2021) in the proof of their Theorem 1. Our constructions differ in the set of target policies $\Pi_{M}\in\Pi$ which are defined in Appendix D.2.
@@ D.2 Instance of the class
See Appendix B for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(x)$. Note that the transition function is the same for all MDPs in the class $\mathcal{M}$ and is defined in Appendix D.1. In particular, the two MDPs $M_{w,+}$ or $M_{w,-}$ only differ in their reward functions, which are opposite.
@@
s.t $\dim B_{K}>0$ and $w\in B_{K}$. Set $B_{K+1}=\langle w\rangle$. Let $H_{w}=\{B_{1},...,B_{K},B_{K+1}\}$ denote the set of nested subspaces (including $\langle w\rangle$). This is not defined as an ordered set, but for notational purposes the order can always be recovered since the sequence must be nested. If $A$ is a subspace of $\mathbb{R}^{d}$, let $\text{proj}_{A}(x)$ denote the orthogonal projection of $x$ onto $A$. See Appendix B for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$ and sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. A target policy $\pi_{H_{w}}$ is specified by the set of nested subspaces $H_{w}$ and is defined as:
@@
  • For $k=0,...,K$, if $s\in\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k})\setminus\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k+1})$, then $s\notin\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k+1})$ and from Definition B.2, this means that for any $u\in B_{k+1}$, $|u^{T}s|/\|u\|_{2}\|s\|_{2}\leq\gamma$. Denoting $x=\text{proj}_{B_{k+1}}(s)\in B_{k+1}$,
@@
  • If $s\in\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{K+1})\setminus(\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w))$, , then $s\notin(\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w))$ and from Definition B.1, this means that $|w^{T}s|\leq\gamma$. Denoting $x=\text{proj}_{B_{K+1}}(s)\in B_{K+1}$,
@@
D.3 Realizability:
We show that the action-value of any target policy for any MDP $M\in\mathcal{M}$ can be linearly represented with the feature map defined in Appendix D.1. The action-value of a target policy $\pi_{H_{w}}$ only depends on $w$ and the sign $\pm$ of the reward function of the MDP, the sequence of nested subspace does not matter beyond $w$.
Lemma D.1 ($Q^{\pi}$ Realizability).
@@ D.3 Realizability:
D.4 Proof of Theorem 4.4
Consider the MDP class described in Appendix D.1 and Appendix D.2. First, we know from Lemma D.1 ($Q^{\pi}$ Realizability) that all instances of the PE problem characterised by the MDP class $\mathcal{M}$ and target policies $\Pi$ satisfy Assumption 4.1 ($Q^{\pi}$ is Realizable) with the feature map $\phi(\cdot,\cdot)$ defined in Appendix D.1.
@@
The learner is sample-efficient so $n$ is polynomial in $d$. Specifically, there exists some constant $\alpha>0$ and some integer $T$ such that $n\leq\alpha d^{T}$. Consider $N\in\mathbb{N}$ s.t $2^{N}\leq d<2^{N+1}$ and set $d_{+}=2^{N}$. Note that $d_{+}\geq d/2$. Fix $W=exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{K}})$. If $n\leq W$, we can show the learner cannot solve the PE problem. Consider both cases:
D.4.1 Case 1
@@
D.4.2 Case 2
$n\leq W\implies n_{k}\leq\exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{k}})$ for $k=1,...,K$. We inductively define the sequence of nested subspaces $B_{K}\subset...\subset B_{1}$ characterising $w$ ($\in B_{K}$) and used by the target policy $\pi_{M}$ s.t for $k=1,...,K$, $\dim B_{k}=2^{\lceil N/4^{k}\rceil}\geq 2^{8}$ and
@@ D.4.2 Case 2
We proceed by induction. Let $B_{K}\subset...\subset B_{1}$ be an arbitrary sequence of subspaces and $w\in B_{K}\cap\mathcal{B}$ arbitrary. We will define these but for now consider the target policy $\pi_{H_{w}}$ (see Appendix D.2) with $H_{w}=\{\langle w\rangle,B_{K},...,B_{1}\}$.
@@
so by Lemma C.3 there exists a subspace $H\in\mathcal{G}_{2^{\lceil N/4\rceil},d_{+}}(\mathbb{R})$ of dimension $2^{\lceil N/4\rceil}$ ($\geq d_{+}^{1/4}$) such that
@@
Inductive Step: Suppose that at round $k$, there exists a sequence of nested subspaces $B_{k}\subset...\subset B_{1}$ used by the target policy $\pi_{H_{w}}$ s.t for $i=1,...,k$, $\dim B_{i}=2^{\lceil N/4^{i}\rceil}\geq 2^{8}$ and
@@
Then we know that the number of those queries that are in $B_{k}$ is also less than $\Big{(}\frac{d_{k}}{2}\Big{)}^{\frac{1}{8}g(\gamma)d_{k}^{1/4}}$. Therefore by Lemma C.3, there exists a subspace $H\subset B_{k}$ of dimension $2^{\lceil\lceil N/4^{k}\rceil/4\rceil}\geq d_{k+1}=2^{\lceil N/4^{k+1}\rceil}$ (can reduce the dimension if they do not match, removing dimensions will not add queries to $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$) such that
@@
We remark Lemma C.3 establishes the existence of a subspace in $\mathcal{G}_{2^{\lceil\lceil N/4^{k}\rceil/4\rceil},d_{k}}(\mathbb{R})$ where the ambient space is $\mathbb{R}^{d_{k}}$ instead of $\mathbb{R}^{d}$. Any $d_{k}$-dimensional subspace of $\mathbb{R}^{d}$ is isomorphic to $\mathbb{R}^{d_{k}}$, so we can consider $B_{k}$, project the points of $A_{k+1}$ into $B_{k}$, apply Lemma C.3 to get $H$ within $B_{k}$, and extend $H$ to be defined in $\mathbb{R}^{d}$. We omit these steps here for clarity but detail them fully in Appendix D.4.3. This also holds for the base case (where $d_{+}$ is used as the ambient dimension instead of $d$)
When applying Lemma C.3 with ambient dimension $d_{k}=2^{\lceil N/4^{k}\rceil}$, we require $\lceil N/4^{k}\rceil\geq 8$. Since $d_{k}\geq d_{+}^{1/4^{k}}$ and $d_{+}\geq d/2$, it is enough to have
@@
The learner has to produce an estimate of $Q^{\pi_{H_{w}}}(\bar{s},\cdot)$ for $\bar{s}=\bm{0}$. But from Lemma D.1,
@@
D.4.3 Dealing with switches in ambient space
In this section we write out the missing details of the proof of Theorem 4.4 in Appendix D.4. In particular, assuming the condition on $n_{k+1}$ is satisfied, we show that we can use Lemma C.3 for the existence of a subspace $H\subset B_{k}\subset\mathbb{R}^{d}$ of dimension $d_{k+1}$ s.t
@@ D.4.3 Dealing with switches in ambient space
Fix $k\in\{0,1,...,K-1\}$ and set $d_{0}=d_{+}$ and $B_{0}$ to any $d_{+}$-dimensional subspace of $\mathbb{R}^{d}$. Picking up the proof of in Appendix D.4, the condition on $n_{k+1}=|A_{k+1}|$ for Lemma C.3 is satisfied. $B_{k}\in\mathcal{G}_{d_{k},d}(\mathbb{R})$ is a $d_{k}$-dimensional subspace within $\mathbb{R}^{d}$. We find an orthonormal basis for $B_{k}$:
@@
The size of $A_{k+1}^{p}$ is $|A_{k+1}^{p}|\leq|A_{k+1}|=n_{k+1}$ so we can apply Lemma C.3: there exists a subspace $H^{p}\in\mathcal{G}_{d_{k+1},d_{k}}(\mathbb{R})$ s.t
@@
Appendix E Proof of Theorem 4.5
E.1 MDP Class

### v2

@@ D.3.1 Preliminary Lemmas
This lemma is similar to Lemma D.4 but uses a more explicit notion of distance between subspaces. The proof is given in Appendix D.3.4.
D.3.2 Proof of Lemma D.3
To prove Lemma D.3, we show the existence of a subspace packing $\mathcal{C}$ in $\mathcal{G}_{m,d}(\mathbb{R})$ (with $m=2^{\lceil N/4\rceil}$) and use Lemma D.4. The subspace packing must satisfy two conditions:
@@
To show the existence of a suitable subspace packing, we use Lemma D.2 with $k=\lfloor\frac{g(\gamma)}{m}\sqrt{d}+1\rfloor$ ($\leq d$). Lemma D.2 gives a packing $\mathcal{C}$ of size $|\mathcal{C}|\geq\Big{(}\frac{d}{2}\Big{)}^{\frac{g(\gamma)}{2m}\sqrt{d}-1}\lfloor\frac{d}{m}\rfloor$ s.t
@@
the packing $\mathcal{C}$ satisfies $|\mathcal{C}|\geq n+1$ and $d_{\min}(\mathcal{C})\geq\sqrt{m-g(\gamma)^{2}}$. By Lemma D.4 there exists a subspace $A\in G_{m,d}(\mathbb{R})$ of dimension $m=2^{\lceil N/4\rceil}$ s.t.
@@
D.3.3 Proof of Lemma D.4
Consider distinct $A,B\in\mathcal{C}$, i.e $A,B$ are subspaces of dimension $m$ s.t.
@@
Now letting $20\leq\theta_{1}\leq...\leq\theta_{m}\leq\pi/$ 2 be the principal angles between $A$ and $B$ (see Appendix D), we have
@@
The first principal angle $\theta_{1}=\arccos a^{T}b$ where $a\in A$, $b\in B$ are unit vectors chosen s.t.
@@
Since this applies to arbitrary distinct $A,B\in\mathcal{C}$, there are $(n+1)$ subspaces $A_{1},...,A_{n+1}\in\mathcal{C}$ of dimension $m$ s.t. for any $i\neq j$,
@@
By Lemma D.5, there exists a subspace $H$ in $\{A_{1},...,A_{n+1}\}$ s.t. for all $y\in\{y_{1},...,y_{n}\}$,
@@
D.3.4 Proof of Lemma D.5
Recall that we consider $(n+1)$ subspaces $A_{1},...,A_{n+1}$ of dimension $m$ s.t. for any $i\neq j$,
@@
Appendix E Proof of Theorem 4.4
E.1 MDP Class
The PE problems used in the proof of Theorem 4.4 are characterised by a class of MDPs $\mathcal{M}$ and target policies $\Pi$. In this section, we define the class of MDPs $\mathcal{M}$. All MDPs $M\in\mathcal{M}$ in the class share the same state-space $\mathcal{S}$, action space $\mathcal{A}$, feature map and transition function $p$ but differ in the reward function $r_{M}$ and target policy $\pi_{M}$. This class of MDPs is the same as the one used by Zanette (2021) in the proof of their Theorem 1. Our constructions differ in the set of target policies $\Pi_{M}\in\Pi$ which are defined in Appendix E.2.
@@ E.2 Instance of the class
See Appendix C for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(x)$. Note that the transition function is the same for all MDPs in the class $\mathcal{M}$ and is defined in Appendix E.1. In particular, the two MDPs $M_{w,+}$ or $M_{w,-}$ only differ in their reward functions, which are opposite.
@@
s.t. $\dim B_{K}>0$ and $w\in B_{K}$. Set $B_{K+1}=\langle w\rangle$. Let $H_{w}=\{B_{1},...,B_{K},B_{K+1}\}$ denote the set of nested subspaces (including $\langle w\rangle$). This is not defined as an ordered set, but for notational purposes the order can always be recovered since the sequence must be nested. If $A$ is a subspace of $\mathbb{R}^{d}$, let $\text{proj}_{A}(x)$ denote the orthogonal projection of $x$ onto $A$. See Appendix C for the definition of hyper-spherical caps $\mathcal{C}_{\gamma}(w)$ and sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$. A target policy $\pi_{H_{w}}$ is specified by the set of nested subspaces $H_{w}$ and is defined as:
@@
  • For $k=0,...,K$, if $s\in\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k})\setminus\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k+1})$, then $s\notin\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{k+1})$ and from Definition C.2, this means that for any $u\in B_{k+1}$, $|u^{T}s|/\|u\|_{2}\|s\|_{2}\leq\gamma$. Denoting $x=\text{proj}_{B_{k+1}}(s)\in B_{k+1}$,
@@
  • If $s\in\overset{\triangle}{\mathcal{C}}_{\gamma}(B_{K+1})\setminus(\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w))$, , then $s\notin(\mathcal{C}_{\gamma}(w)\cup\mathcal{C}_{\gamma}(-w))$ and from Definition C.1, this means that $|w^{T}s|\leq\gamma$. Denoting $x=\text{proj}_{B_{K+1}}(s)\in B_{K+1}$,
@@
E.3 Realizability:
We show that the action-value of any target policy for any MDP $M\in\mathcal{M}$ can be linearly represented with the feature map defined in Appendix E.1. The action-value of a target policy $\pi_{H_{w}}$ only depends on $w$ and the sign $\pm$ of the reward function of the MDP, the sequence of nested subspace does not matter beyond $w$.
Lemma E.1 ($Q^{\pi}$ Realizability).
@@ E.3 Realizability:
E.4 Proof of Theorem 4.4
Consider the MDP class described in Appendix E.1 and Appendix E.2. First, we know from Lemma E.1 ($Q^{\pi}$ Realizability) that all instances of the PE problem characterised by the MDP class $\mathcal{M}$ and target policies $\Pi$ satisfy Assumption 4.1 ($Q^{\pi}$ is Realizable) with the feature map $\phi(\cdot,\cdot)$ defined in Appendix E.1.
@@
The learner is sample-efficient so $n$ is polynomial in $d$. Specifically, there exists some constant $\alpha>0$ and some integer $T$ such that $n\leq\alpha d^{T}$. Consider $N\in\mathbb{N}$ s.t. $2^{N}\leq d<2^{N+1}$ and set $d_{+}=2^{N}$. Note that $d_{+}\geq d/2$. Fix $W=\exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{K}})$. If $n\leq W$, we can show the learner cannot solve the PE problem. Consider both cases:
E.4.1 Case 1
@@
E.4.2 Case 2
$n\leq W\implies n_{k}\leq\exp(\frac{1}{8}g(\gamma)d_{+}^{1/4^{k}})$ for $k=1,...,K$. We inductively define the sequence of nested subspaces $B_{K}\subset...\subset B_{1}$ characterising $w$ ($\in B_{K}$) and used by the target policy $\pi_{M}$ s.t. for $k=1,...,K$, $\dim B_{k}=2^{\lceil N/4^{k}\rceil}\geq 2^{8}$ and
@@ E.4.2 Case 2
We proceed by induction. Let $B_{K}\subset...\subset B_{1}$ be an arbitrary sequence of subspaces and $w\in B_{K}\cap\mathcal{B}$ arbitrary. We will define these but for now consider the target policy $\pi_{H_{w}}$ (see Appendix E.2) with $H_{w}=\{\langle w\rangle,B_{K},...,B_{1}\}$.
@@
so by Lemma D.3 there exists a subspace $H\in\mathcal{G}_{2^{\lceil N/4\rceil},d_{+}}(\mathbb{R})$ of dimension $2^{\lceil N/4\rceil}$ ($\geq d_{+}^{1/4}$) such that
@@
Inductive Step: Suppose that at round $k$, there exists a sequence of nested subspaces $B_{k}\subset...\subset B_{1}$ used by the target policy $\pi_{H_{w}}$ s.t. for $i=1,...,k$, $\dim B_{i}=2^{\lceil N/4^{i}\rceil}\geq 2^{8}$ and
@@
Then we know that the number of those queries that are in $B_{k}$ is also less than $\Big{(}\frac{d_{k}}{2}\Big{)}^{\frac{1}{8}g(\gamma)d_{k}^{1/4}}$. Therefore by Lemma D.3, there exists a subspace $H\subset B_{k}$ of dimension $2^{\lceil\lceil N/4^{k}\rceil/4\rceil}\geq d_{k+1}=2^{\lceil N/4^{k+1}\rceil}$ (can reduce the dimension if they do not match, removing dimensions will not add queries to $\overset{\triangle}{\mathcal{C}}_{\gamma}(H)$) such that
@@
We remark Lemma D.3 establishes the existence of a subspace in $\mathcal{G}_{2^{\lceil\lceil N/4^{k}\rceil/4\rceil},d_{k}}(\mathbb{R})$ where the ambient space is $\mathbb{R}^{d_{k}}$ instead of $\mathbb{R}^{d}$. Any $d_{k}$-dimensional subspace of $\mathbb{R}^{d}$ is isomorphic to $\mathbb{R}^{d_{k}}$, so we can consider $B_{k}$, project the points of $A_{k+1}$ into $B_{k}$, apply Lemma D.3 to get $H$ within $B_{k}$, and extend $H$ to be defined in $\mathbb{R}^{d}$. We omit these steps here for clarity but detail them fully in Appendix E.4.3. This also holds for the base case (where $d_{+}$ is used as the ambient dimension instead of $d$)
When applying Lemma D.3 with ambient dimension $d_{k}=2^{\lceil N/4^{k}\rceil}$, we require $\lceil N/4^{k}\rceil\geq 8$. Since $d_{k}\geq d_{+}^{1/4^{k}}$ and $d_{+}\geq d/2$, it is enough to have
@@
The learner has to produce an estimate of $Q^{\pi_{H_{w}}}(\bar{s},\cdot)$ for $\bar{s}=\bm{0}$. But from Lemma E.1,
@@
E.4.3 Dealing with switches in ambient space
In this section we write out the missing details of the proof of Theorem 4.4 in Appendix E.4. In particular, assuming the condition on $n_{k+1}$ is satisfied, we show that we can use Lemma D.3 for the existence of a subspace $H\subset B_{k}\subset\mathbb{R}^{d}$ of dimension $d_{k+1}$ s.t.
@@ E.4.3 Dealing with switches in ambient space
Fix $k\in\{0,1,...,K-1\}$ and set $d_{0}=d_{+}$ and $B_{0}$ to any $d_{+}$-dimensional subspace of $\mathbb{R}^{d}$. Picking up the proof of in Appendix E.4, the condition on $n_{k+1}=|A_{k+1}|$ for Lemma D.3 is satisfied. $B_{k}\in\mathcal{G}_{d_{k},d}(\mathbb{R})$ is a $d_{k}$-dimensional subspace within $\mathbb{R}^{d}$. We find an orthonormal basis for $B_{k}$:
@@
The size of $A_{k+1}^{p}$ is $|A_{k+1}^{p}|\leq|A_{k+1}|=n_{k+1}$ so we can apply Lemma D.3: there exists a subspace $H^{p}\in\mathcal{G}_{d_{k+1},d_{k}}(\mathbb{R})$ s.t.
@@
E.5 Illustration of MDP construction
In this section, we provide a high-level illustration of how the hard instance $M\in\mathcal{M}$ is constructed when $d=3$ and the learner makes three queries per round. The reward vector of the MDP is 00 everywhere except in the $\gamma$-hyper-spherical sector $\overset{\triangle}{\mathcal{C}}_{\gamma}(w)$ (see Appendix C) of a vector $w\in\mathcal{B}$, which is unknown to the learner. The target policy will be constructed in such a way that $Q^{\pi}(s,a)=\pm\phi(s,a)^{T}w$. If the learner can identify $w$ and make a query in its $\gamma$ hyper-spherical sector (to distinguish between the $+$ and the --), then it can fully solve the PE problem. Therefore the aim of the learner in choosing its queries is to identify this vector $w$.
Figure 3: Round 1.
Round 1 (Figure 3).
  • A: Three queries $\phi_{i}=\phi(s_{i},a_{i})=a_{i}$ ($i=1,2,3$) are made by the learner.
  • B: The three queries with their hyper-spherical sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(\phi_{i})$ ($i=1,2,3$) (we only show one side of the sector for clarity).
  • C: There exists a $2$-dimensional subspace $B_{1}$ that does not intersect the hyper-spherical sectors of the queries which the environment can use to further hide $w$ if $B_{1}$ does not intersect the $\gamma$ hyper-spherical sectors of the queries, then the queries $\phi_{i}$ will not be in the $\gamma$ hyper-spherical sector of any point in $B_{1}$ (in particular $w$ and cannot identify the $+$ or -- in the reward).
  • D: The target-policy $\pi_{M}$ is constructed in such a way that realizability is satisfied with $w\in B_{1}$. In particular, $\pi_{M}(s_{i}^{+})=\frac{1}{\gamma}\text{proj}_{B_{1}}(s_{i}^{+})=\frac{1}{\gamma}\text{proj}_{B_{1}}(a_{i})$ is valid because $B_{1}$ does not intersect the sectors of the queries (see Appendix E.2). This choice allows the component of $\phi$ in the direction of $w$ to grow by $1/\gamma$ in $\phi^{+}$, this will cancel with the $\gamma$ discount in the bellman equation and is a key step in satisfying realizability.
    Since the orthogonal projection is the same for all vectors in $B_{1}$, this definition of the target-policy (allowing realizability to be satisfied - see Appendix E.3) does not reveal any additional information about $w$ other than $w\in B_{1}$. In addition, $w$ or the actions taken by the policy within $B_{1}$ do not need to be specified yet and can depend on queries made by the learner in later rounds.
    Since $\phi_{i}–\gamma\phi^{+}_{i}=a_{i}-\text{proj}_{B_{1}}(a_{i})$ is orthogonal to $B_{1}$, this construction admits the interpretation of erasing information along the directions of the subspace $B_{1}$. Hence, this provides the intuition discussed in Section 5 but is in fact exactly what is required to hide the $w$ vector within the subspace $B_{1}$.
Figure 4: Round 2.
Round 2 (Figure 4). Without loss of generality, we assume the queries made by the learner are within $B_{1}$ (the MDP is fully identified outside of $B_{1}$ so the learner can gain no further information from queries outside of $B_{1}$) and only visualise $B_{1}$ as a circle embedded in the sphere from Figure 3.
  • E: Three queries $\phi_{i}=\phi(s_{i},a_{i})=a_{i}$ ($i=4,5,6$) are made by the learner.
  • F: The three queries with their ($2$-dimensional) hyper-spherical sectors $\overset{\triangle}{\mathcal{C}}_{\gamma}(\phi_{i})$ ($i=4,5,6$).
  • G: There exists a $1$-dimensional subspace $B_{2}=\langle w\rangle$ that does not intersect the hyper-spherical sectors of the queries. As in C above, if $\langle w\rangle$ does not intersect the $\gamma$ hyper-spherical sectors of the queries, then the queries $\phi_{i}$ will not be in the $\gamma$ hyper-spherical sector of $w$.
  • H: Similar to D, the target-policy $\pi_{M}$ is constructed within $B_{1}$ in such a way that realizability is satisfied. In particular, $\pi_{M}(s_{i}^{+})=\frac{1}{\gamma}\text{proj}_{B_{2}}(s_{i}^{+})=\frac{1}{\gamma}(w^{T}a_{i})w$ is valid because $B_{2}=\langle w\rangle$ does not intersect the sectors of the queries. The reward inside $\overset{\triangle}{\mathcal{C}}_{\gamma}(w)$ is characterised by a sign $+$ or --. No queries have been made inside $\overset{\triangle}{\mathcal{C}}_{\gamma}(w)$ and so a third round is required to identify the sign $+$ or --.
This illustration uses three queries per round. In the proof, we formally show that if the number of queries in the first round is polynomial in $d$, then there always exists a subspace of dimension $d^{1/4}$ that does not intersect the sectors of the queries in the first round. We make a similar argument for later rounds to show $\log\log d$ rounds are required to identify the vector $w$. See the earlier sections for all the details.
Appendix F Proof of Theorem 4.5
F.1 MDP Class